# Orchestrator-Agent Trust: A Modular Agentic AI Visual Classification System with Trust-Aware Orchestration and RAG-Based Reasoning

### v2

@@ Abstract
Modern Artificial Intelligence (AI) increasingly relies on multi-agent architectures that blend visual and language understanding. Yet, a pressing challenge remains: How can we trust these agents especially in zero-shot settings with no fine-tuning? We introduce a novel modular Agentic AI visual classification framework that integrates generalist multimodal agents with a non-visual reasoning orchestrator and a Retrieval-Augmented Generation (RAG) module. Applied to apple leaf disease diagnosis, we benchmark three configurations: (I) zero-shot with confidence-based orchestration, (II) fine-tuned agents with improved performance, and (III) trust-calibrated orchestration enhanced by CLIP-based image retrieval and re-evaluation loops. Using confidence calibration metrics (ECE, OCR, CCC), the orchestrator modulates trust across agents. Our results demonstrate a 77.94% accuracy improvement in the zero-shot setting using trust-aware orchestration and RAG, achieving 85.63% overall. GPT-4o showed better calibration, while Qwen-2.5-VL displayed overconfidence. Furthermore, image-RAG grounded predictions with visually similar cases, enabling correction of agent overconfidence via iterative re-evaluation. The proposed system separates perception (vision agents) from meta-reasoning (orchestrator), enabling scalable and interpretable multi-agent AI. This blueprint is extensible to diagnostics, biology, and other trust-critical domains. All models, prompts, results, and system components including the complete software source code are openly released to support reproducibility, transparency, and community benchmarking at our Github page
@@
agentic ai, orchestrator agent trust, trust orchestration, visual classification, retrieval augmented reasoning
@@ 1 Introduction
The integration of vision and language models into autonomous decision-making systems has redefined the boundaries of artificial intelligence (AI), especially in fields that demand both perceptual accuracy and interpretability [1, 2]. Multimodal large language models (LMMs), capable of reasoning over both visual and textual data, are increasingly employed in diverse domains, ranging from autonomous robotics and medical imaging to scientific diagnostics and agricultural monitoring [3, 4, 5]. However, as these models are deployed in high-stakes environments, a fundamental challenge persists: can these systems be trusted to make reliable, transparent, and justifiable decisions particularly in zero-shot or open-world scenarios where no prior task-specific fine-tuning is possible?
@@
Recent developments in agentic AI systems where autonomous agents collaborate, reason, and interact with their environment have highlighted the importance of meta-reasoning and modularity in AI architectures [6, 7, 8]. Rather than relying on monolithic end-to-end networks, agentic systems distribute cognitive tasks across specialized agents [9]. In vision-language applications, this paradigm enables distinct agents to independently process visual inputs, generate explanations, and assess confidence, while an external orchestrator or supervisor performs higher-order reasoning to synthesize their outputs [10, 11, 12]. This architectural shift not only mirrors aspects of human collaborative problem-solving but also introduces an additional layer of oversight, which is critical for ensuring accountability and trust [13].
@@
Despite the promise of this approach, trust calibration within agentic AI remains an underexplored area. Conventional ensemble systems often assume that agents’ self-reported confidence scores are reliable proxies for correctness [14, 15, 16, 17], an assumption that breaks down in zero-shot generalist models such as GPT-4o or Qwen-2.5-VL. These models may exhibit systematic overconfidence or fail to discriminate between subtle categories in domain-specific tasks [18]. Consequently, there is a pressing need to assess and correct misalignments between an agent’s confidence and its actual performance. This problem is particularly acute in scientific and diagnostic domains, where misclassification may lead to erroneous conclusions or interventions.
@@
In visual classification tasks such as plant disease detection, trustworthiness is not merely a matter of predictive accuracy but of interpretability and justification [19, 20, 21]. For instance, two models may arrive at the same prediction but differ significantly in the rationale behind their decision [22, 23]. A reliable AI system should not only be accurate but also capable of articulating why a decision was made and when to defer to alternative evidence [24]. RAG approaches where models consult external databases or prior examples to refine their outputs offer a promising pathway to ground predictions in visual context, enhancing both trust and transparency [25, 26, 27].
@@
To address these limitations, we propose a novel Agentic AI framework for visual classification that integrates trust-aware orchestration with retrieval-augmented reasoning in a modular, interpretable, and zero-shot-capable architecture. Our system coordinates multiple multimodal agents GPT-4o and Qwen-2.5-VL with a non-visual orchestrator that synthesizes final predictions based on reported confidence, natural language justifications, and internal trust scores. When discrepancies or low-confidence predictions arise, the orchestrator triggers a re-evaluation loop using RAG supported by CLIP-based image retrieval, allowing the system to reflect and refine its decision using similar visual precedents. We introduce a three-stage pipeline (see schematic box below), incrementally augmenting the system’s capabilities:
[…]
@@
  • Experiment I: Agents operate in zero-shot mode; the orchestrator selects final predictions using reported confidence scores.
  • Experiment II: Agents are fine-tuned on a curated apple leaf disease dataset using hyperparameter optimization (Bayesian search) to evaluate the effects of supervised domain adaptation.
@@ 2.1 Zero-Shot Agentic Classification with Confidence-Aware Orchestration
Figure 2b reports top-1 classification accuracy for each agent, where GPT-4o clearly outperforms the others. Weighted precision scores are illustrated in Fig.2c, and raw confidence distributions are presented in Fig.2d. Average confidence levels per model are plotted in Fig.2e, with Qwen showing systematic overconfidence. Recall and F1 scores are detailed in Figs.2f and  2g, respectively, showing GPT-4o leading across metrics.
Confusion matrices in Figs. 2h,  2i, and  2j display inter-class prediction patterns for Qwen, GPT, and the orchestrator, respectively. Misclassification frequently occurred between visually similar disease classes (e.g., black-rot and rust), highlighting the inherent difficulty of this task for zero-shot models.
@@ II. Few-Shot + Confidence-Aware
Accuracy: 0.9813 Precisionw: 0.9817 Recallw: 0.9813 F1-scorew: 0.9812 Precisionm: 0.9817 Recallm: 0.9813 F1-scorem: 0.9812
@@ 2.2 Fine-Tuned Agentic Models with Confidence-Aware Orchestration
To evaluate whether supervised adaptation improves model reliability and ensemble synergy, we conducted a second experiment in which both GPT-4o and Qwen-2.5-VL were fine-tuned on the apple disease classification dataset. The overall setup mirrored Experiment 2.1, retaining identical multimodal inputs, test prompts, and classification objectives. However, the zero-shot agents were replaced with domain-adapted variants, each fine-tuned on a labeled training set of 512 samples and validated on 128 additional samples. Importantly, the orchestrator (o3-mini-2025-01-31) remained unchanged and continued to perform confidence-aware arbitration, relying solely on agent-generated confidence scores and textual justifications to make final decisions.
@@
Recall and F1 improvements are presented in Figs. 3f and 3g, respectively. The F1 score for GPT-4o increased from 50.78% to 98.12%, and for Qwen from 37.63% to 95.58%. These gains reveal that fine-tuning not only enhances correct predictions but also reduces false positives and negatives. Confusion matrices in Figs. 3h–3j show sharply improved class-level discrimination. Notably, categories like black-rot and rust, which previously showed misclassification overlap, now exhibit strong diagonal dominance.
@@
Detailed fine-tuning configurations for each agent model are presented in Section 4. The results confirm that supervised adaptation significantly enhances the classification performance of general-purpose vision-language agents within the agentic framework. After fine-tuning on the domain-specific apple disease dataset, both GPT-4o and Qwen-2.5-VL showed notable gains across accuracy, precision, recall, and F1 scores. As shown in Table 1, GPT-4o achieved 98.13% accuracy and Qwen 95.63%, while the orchestrator despite lacking direct image access reached 97.50% accuracy. This highlights how confidence-aware arbitration benefits from well-calibrated agent responses, with the orchestrator’s performance reflecting the value of alignment between self-reported confidence and correctness.
@@ 2.3 Trust-Aware Orchestration with RAG and Re-Evaluation Loops
[…]
Performance: As summarized in Table 1, GPT-4o achieved 87.50% accuracy and Qwen reached 73.13%, both significantly outperforming their initial zero-shot baselines in Experiment I. The orchestrator attained an accuracy of 85.63%, demonstrating that trust-aware orchestration, combined with retrieval-based contextualization, can close a large portion of the performance gap typically addressed via fine-tuning. Subfigure 4b shows the accuracy improvements, while subfigures 4c, 4f, and 4g illustrate the precision, recall, and F1-score gains, respectively.
@@
Trust profiling: From the profiling, several insights emerge regarding agent reliability under zero-shot conditions. Importantly, both models exhibit suboptimal calibration and trustworthiness metrics overall (Table 1) an expected outcome given that they were not fine-tuned on the task-specific dataset. Nonetheless, relative differences offer valuable guidance for orchestration design:
@@
Collectively, Experiment III highlights that zero-shot generalist agents, when embedded in trust-calibrated agentic systems with access to external retrieval, can achieve expert-level accuracy in image classification tasks. These findings support a broader vision of scalable agentic intelligence where trust, not tuning, becomes the key to real-world deployment.
@@
Comparative Performance and Ablation Analysis
Time and Calibration Performance Across Configurations.

### v3

@@ Abstract
Modern Artificial Intelligence (AI) increasingly relies on multi-agent architectures that blend visual and language understanding. Yet, a pressing challenge remains: How can we trust these agents especially in zero-shot settings with no fine-tuning? We introduce a novel modular Agentic AI visual classification framework that integrates generalist multimodal agents with a non-visual reasoning orchestrator and a Retrieval-Augmented Generation (RAG) module. Applied to apple leaf disease diagnosis, we benchmark three configurations: (I) zero-shot with confidence-based orchestration, (II) fine-tuned agents with improved performance, and (III) trust-calibrated orchestration enhanced by CLIP-based image retrieval and re-evaluation loops. Using confidence calibration metrics (ECE, OCR, CCC), the orchestrator modulates trust across agents. Our results demonstrate a 77.94% accuracy improvement in the zero-shot setting using trust-aware orchestration and RAG, achieving 85.63% overall. GPT-4o showed better calibration, while Qwen-2.5-VL displayed overconfidence. Furthermore, image-RAG grounded predictions with visually similar cases, enabling correction of agent overconfidence via iterative re-evaluation. The proposed system separates perception (vision agents) from meta-reasoning (orchestrator), enabling scalable and interpretable multi-agent AI. This blueprint illustrates how Agentic AI can deliver trustworthy, modular, and transparent reasoning, and is extensible to diagnostics, biology, and other trust-critical domains. In doing so, we highlight Agentic AI not just as an architecture but as a paradigm for building reliable multi-agent intelligence. All models, prompts, results, and system components including the complete software source code are openly released to support reproducibility, transparency, and community benchmarking at our Github page.
@@
[…]
@@ 1 Introduction
The integration of vision and language models into autonomous decision-making systems has redefined the boundaries of artificial intelligence (AI), especially in fields that demand both perceptual accuracy and interpretability Afroogh et al. (2024); Ilievski (2025). Multimodal large language models (LMMs), capable of reasoning over both visual and textual data, are increasingly employed in diverse domains, ranging from autonomous robotics and medical imaging to scientific diagnostics and agricultural monitoring Bradshaw et al. (2025); Yang et al. (2024); Mon-Williams et al. (2025). However, as these models are deployed in high-stakes environments, a fundamental challenge persists: can these systems be trusted to make reliable, transparent, and justifiable decisions particularly in zero-shot or open-world scenarios where no prior task-specific fine-tuning is possible?
@@
Recent developments in agentic AI systems where autonomous agents collaborate, reason, and interact with their environment have highlighted the importance of meta-reasoning and modularity in AI architectures Ale et al. (2025); Buehler (2025); Sapkota et al. (2026). Rather than relying on monolithic end-to-end networks, agentic AI systems distribute cognitive tasks across specialized AI agents Savaglio et al. (2020). In vision-language applications, this paradigm enables distinct agents to independently process visual inputs, generate explanations, and assess confidence, while an external orchestrator or supervisor performs higher-order reasoning to synthesize their outputs Jeyakumar et al. (2024); Zhai et al. (2024); Lin et al. (2025). This architectural shift not only mirrors aspects of human collaborative problem-solving but also introduces an additional layer of oversight, which is critical for ensuring accountability and trust Qiao et al. (2025).
@@
Despite the promise of this approach, trust calibration within agentic AI remains an underexplored area. Conventional ensemble systems often assume that agents’ self-reported confidence scores are reliable proxies for correctness Ma et al. (2024b); Warmsley et al. (2025); Wang et al. (2025); Huang et al. (2024), an assumption that breaks down in zero-shot generalist models such as GPT-4o or Qwen-2.5-VL. These models may exhibit systematic overconfidence or fail to discriminate between subtle categories in domain-specific tasks Wen et al. (2024). Consequently, there is a pressing need to assess and correct misalignments between an agent’s confidence and its actual performance. This problem is particularly acute in scientific and diagnostic domains, where misclassification may lead to erroneous conclusions or interventions.
@@
In visual classification tasks such as plant disease detection, trustworthiness is not merely a matter of predictive accuracy but of interpretability and justification Nigar et al. (2024); Ding et al. (2024); Qadri et al. (2024). For instance, two models may arrive at the same prediction but differ significantly in the rationale behind their decision Ali et al. (2023); Bajorath (2025). A reliable AI system should not only be accurate but also capable of articulating why a decision was made and when to defer to alternative evidence Messeri and Crockett (2024). RAG approaches where models consult external databases or prior examples to refine their outputs offer a promising pathway to ground predictions in visual context, enhancing both trust and transparency Ke et al. (2025); Dong et al. (2025); Tozuka et al. (2025).
@@
To address these limitations, we propose a novel Agentic AI framework for visual classification that integrates trust-aware orchestration with retrieval-augmented reasoning in a modular, interpretable, and zero-shot-capable architecture. Our system coordinates multiple multimodal agents GPT-4o and Qwen-2.5-VL with a non-visual orchestrator that synthesizes final predictions based on reported confidence, natural language justifications, and internal trust scores. When discrepancies or low-confidence predictions arise, the orchestrator triggers a re-evaluation loop using RAG supported by CLIP-based image retrieval, allowing the system to reflect and refine its decision using similar visual precedents. We introduce a three-stage pipeline (see Figure 1), incrementally augmenting the system’s capabilities.
Three-Stage Trust-Aware Agentic AI Framework
Experiment IZero-Shot Agents+ Confidence-AwareOrchestrationExperiment IIFine-Tuned Agents+ Confidence-AwareOrchestrationExperiment IIIZero-Shot Agents+ Trust-Aware Orchestration+ RAG + Re-Evaluation LoopAgent Fine-TuningAdd Trust Metrics + RAGGeneralist Multimodal ModelsDomain-Specialized via Supervised Fine-TuningTrust-Calibrated + Evidence-Grounded PipelineGPT-4o, Qwen-2.5-VL (Zero-Shot)GPT-4o, Qwen-2.5-VL (Fine-Tuned)GPT-4o, Qwen-2.5-VL + CLIP (RAG)
Figure 1: Three-stage trust-aware agentic AI framework across experiments.
@@
  • Experiment I: AI Agents operate in zero-shot mode; the orchestrator selects final predictions using reported confidence scores.
  • Experiment II: AI Agents are fine-tuned on a curated apple leaf disease dataset using hyperparameter optimization (Bayesian search) to evaluate the effects of supervised domain adaptation.
@@
2 Methodology
The system comprises three experimental setups.
2.1 Experiment I
Each agent receives a prompt containing a single input image (from the test set) and is tasked with predicting the most appropriate plant disease class. In addition to the predicted class label, agents are required to return (i) a confidence score in the range [0.0–1.0] and (ii) a natural language explanation justifying their decision (Appendix, Fig. A1). These three outputs classification, confidence, and rationale are compiled and forwarded to a non-visual orchestrator model (o3-mini-2025-01-31 OpenAI (2025)), which serves as a comparative reasoner. The orchestrator does not process images directly; instead, it evaluates the agents’ predictions and justifications in light of their associated confidence scores and produces a final classification decision through structured, confidence-aware reasoning (Appendix, Fig. A2).
2.2 Experiment II
  1. 1.
    1. (a)
    2. (b)
  2. 2.
    Qwen-2.5-VL-7B Fine-Tuning: To adapt the Qwen-2.5 Vision-Language 7B (VL-7B) model for our task, we fine-tuned the unsloth/Qwen2.5-VL-7B-Instruct-bnb-4bit model Wang et al. (2024), a 4-bit quantized variant of the original pretrained Qwen2.5 model. This quantized version enables efficient loading and inference with reduced memory requirements, making it suitable for commodity GPU hardware. We employed Low-Rank Adaptation (LoRA) in 16-bit precision to inject trainable adapters into the model, thereby allowing effective fine-tuning without full dequantization of the backbone weights.
2.3 Experiment III
Trust Evaluation Layer. The Trust Evaluation Layer conducts offline trust profiling of the AI agents by quantifying their calibration, discriminative reliability, and consistency over a labeled dataset. Specifically, both agents were evaluated in a zero-shot inference setting on the training image set originally employed in Experiment II for fine-tuning, but here presented without ground-truth labels during evaluation. For each image, we recorded the predicted class, associated confidence score, and explanatory output provided by the agent. Using this comprehensive log of model behavior, we computed a suite of quantitative trustworthiness metrics to assess each agent’s predictive confidence alignment and overall reliability under zero-shot conditions.
  • Overconfidence Ratio (OCR): The OCR measures the proportion of incorrect predictions made with high confidence (e.g., confidence > 0.9), capturing the extent to which models exhibit unjustified certainty. A high OCR value signals a critical misalignment between model confidence and actual correctness, often leading to erroneous decisions in confidence-based orchestration pipelines. In our trust-aware framework, OCR serves as a penalizing factor in agent arbitration, allowing the orchestrator to detect and downweight predictions from agents that frequently make confidently wrong decisions.
    Measuring how often models are confidently wrong.
  • Consistency Gap (CG): The CG quantifies the divergence in a model’s predictions when exposed to semantically equivalent but syntactically varied prompts, serving as a proxy for reasoning stability under linguistic perturbations. A large CG indicates that the AI agent’s output is highly sensitive to prompt phrasing, undermining reliability and reproducibility in decision-making. Within our framework, CG is used as a trust signal to evaluate the robustness of agent responses, enabling the orchestrator to identify models prone to inconsistency and initiate re-evaluation when necessary.
    $\text{CG}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[P_{i}^{(1)}\neq P_{i}^{(2)}],$
    Assessing variation in outputs across prompt formulations.
$\mathbf{e}_{i}=\text{L2\_normalize}(\text{ViT}_{\text{CLIP}}(\mathbf{I}_{i})),\\ \quad\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})=\cos(\theta_{ij})=\mathbf{e}_{i}\cdot\mathbf{e}_{j}$ (1)
All reference embeddings are stored in a vector database implemented using Facebook AI Similarity Search (FAISS), employing an exact inner product index (IndexFlatIP). Each entry is paired with a category label $y_{i}$ taking values healthy, black-rot, rust, scab, and associated metadata (e.g., image URLs, index references).
Each entry is paired with a category label $y_{i}\in\{\text{healthy},\allowbreak\ \text{black-rot},\allowbreak\ \text{rust},\allowbreak\ \text{scab}\}$ and associated metadata (e.g., image URLs, index references).
$R_{k}=\arg\max_{k}\{\cos(\mathbf{e}_{q},\mathbf{e}_{i})\mid i\in[1,N]\}$ (2)
$\text{conf}(c)=\frac{\sum_{i\in R_{k},y_{i}=c}s_{i}}{\sum_{j\in R_{k}}s_{j}}$ (3)
  {"category": "scab", "confidence": 0.5005},
  {"category": "healthy", "confidence": 0.3996},
  {"category": "rust", "confidence": 0.0999}
@@ 3.1 Zero-Shot Agentic Classification with Confidence-Aware Orchestration
Figure 3b reports top-1 classification accuracy for each agent, where GPT-4o clearly outperforms the others. Weighted precision scores are illustrated in Fig.3c, and raw confidence distributions are presented in Fig.3d. Average confidence levels per model are plotted in Fig.3e, with Qwen showing systematic overconfidence. Recall and F1 scores are detailed in Figs.3f and 3g, respectively, showing GPT-4o leading across metrics.
Confusion matrices in Figs. 3h,  3i, and  3j display inter-class prediction patterns for Qwen, GPT, and the orchestrator, respectively. Misclassification frequently occurred between visually similar disease classes (e.g., black-rot and rust), highlighting the inherent difficulty of this task for zero-shot models.
@@ II. Few-Shot + Confidence-Aware
Accuracy: 0.9813
Precisionw: 0.9817
Recallw: 0.9813
F1-scorew: 0.9812
Precisionm: 0.9817
Recallm: 0.9813
F1-scorem: 0.9812
@@ 3.2 Fine-Tuned Agentic Models with Confidence-Aware Orchestration
To evaluate whether supervised adaptation improves model reliability and ensemble synergy, we conducted a second experiment in which both GPT-4o and Qwen-2.5-VL were fine-tuned on the apple disease classification dataset. The overall setup mirrored Experiment I, retaining identical multimodal inputs, test prompts, and classification objectives. However, the zero-shot agents were replaced with domain-adapted variants, each fine-tuned on a labeled training set of 512 samples and validated on 128 additional samples. Importantly, the orchestrator (o3-mini-2025-01-31) remained unchanged and continued to perform confidence-aware arbitration, relying solely on agent-generated confidence scores and textual justifications to make final decisions.
@@
Recall and F1 improvements are presented in Figs. 4f and 4g, respectively. The F1 score for GPT-4o increased from 50.78% to 98.12%, and for Qwen from 37.63% to 95.58%. These gains reveal that fine-tuning not only enhances correct predictions but also reduces false positives and negatives. Confusion matrices in Figs. 4h–4j show sharply improved class-level discrimination. Notably, categories like black-rot and rust, which previously showed misclassification overlap, now exhibit strong diagonal dominance.
@@
The results confirm that supervised adaptation significantly enhances the classification performance of general-purpose vision-language agents within the agentic AI framework. After fine-tuning on the domain-specific apple disease dataset, both GPT-4o and Qwen-2.5-VL showed notable gains across accuracy, precision, recall, and F1 scores. As shown in Table 1, GPT-4o achieved 98.13% accuracy and Qwen 95.63%, while the orchestrator despite lacking direct image access reached 97.50% accuracy. This highlights how confidence-aware arbitration benefits from well-calibrated agent responses, with the orchestrator’s performance reflecting the value of alignment between self-reported confidence and correctness.
@@ 3.3 Trust-Aware Orchestration with RAG and Re-Evaluation Loops
In this setup, low-trust predictions trigger the orchestrator to initiate a re-evaluation loop, prompting agents with their prior responses and augmenting the input with semantic retrievals from Image-RAG. The retrieval module uses CLIP (ViT-B/32) to encode a curated set of class-representative reference images into 512-dimensional embeddings, which are $L_{2}$-normalized and stored in a FAISS vector database using an exact inner product index (IndexFlatIP). Each reference embedding is associated with a disease category label $y_{i}\in{\text{healthy},\text{black-rot},\text{rust},\text{scab}}$ and enriched with metadata such as image URLs and textual definitions. During re-evaluation, the agents receive the top-$k$ most similar reference examples based on cosine similarity, integrated into an updated prompt. This guides them to revise or reaffirm their prediction with additional visual and semantic context. The orchestrator then assesses the updated responses using trust metrics (ECE, OCR, CG) and finalizes the decision based on the most coherent and trustworthy agent output.
Performance: As summarized in Table 1, GPT-4o achieved 87.50% accuracy and Qwen reached 73.13%, both significantly outperforming their initial zero-shot baselines in Experiment I. The orchestrator attained an accuracy of 85.63%, demonstrating that trust-aware orchestration, combined with retrieval-based contextualization, can close a large portion of the performance gap typically addressed via fine-tuning. Subfigure 5b shows the accuracy improvements, while subfigures 5c, 5f, and 5g illustrate the precision, recall, and F1-score gains, respectively.
@@
Trust profiling: From the profiling, several insights emerge regarding agent reliability under zero-shot conditions. Importantly, both models exhibit suboptimal calibration and trustworthiness metrics overall (Table A4) an expected outcome given that they were not fine-tuned on the task-specific dataset. Nonetheless, relative differences offer valuable guidance for orchestration design:
@@
Collectively, Experiment III highlights that zero-shot generalist agents, when embedded in trust-calibrated agentic AI systems with access to external retrieval, can achieve expert-level accuracy in image classification tasks. These findings support a broader vision of scalable agentic intelligence where trust, not tuning, becomes the key to real-world deployment.
@@
3.4 Comparative Performance and Ablation Analysis
3.4.1 Time and Calibration Performance Across Configurations